# Make Me Happier: Evoking Emotions Through Image Diffusion Models

### v3

@@ 4 Experiment
We selected 504 images randomly from the EmoSet dataset Yang et al. (2023), ensuring they were distinct from those in our EmoPair dataset. According to the emotion labels of EmoSet, there are 63 images for each source emotion category. Subsequently, we evaluated both competitive baselines and our EmoEditor on these source images and presented the results in the following subsections.
@@
Ideally, with source images and their classified emotions, our goal is to generate images evoking the other 7 target emotions. Considering emotional valence, we divided the test set into two subsets: within-valence emotion transfer and cross-valence emotion transfer. This division serves two purposes. Firstly, cross-valence emotion transfer is more common in real-world scenarios, like enhancing positive emotions in stressful environments or creating fear in lively settings during Halloween. Secondly, images of the same valence may contain subtle differences challenging for humans to discern. For instance, an amusement park scene might evoke a mixture of excitement and amusement.
@@
Baselines. We compare our EmoEditor with five state-of-the-art methods: (1) Color-transferPitié et al. (2007) (CT); (2) Neural-Style-TransferGatys et al. (2015) (NST). Both methods require an additional reference image from the target emotion category as inputs. We randomly selected these reference images from the EmoSet dataset. We also include representative text-to-image editing models (3) CLIP-StylerKwon, Ye (2022) (Csty) and (4) Ip2p Brooks et al. (2023). We provide these models with the target emotion category as a one-word text prompt, such as “excitement". Emotion-evoked image generation involves three key steps: image understanding, instruction generation, and instruction-based image editing. To tackle these, we concatenate existing large language and vision models and introduce the baseline (5) “Large Model Series" (LMS). This includes BLIP Li et al. (2022) for image captioning, followed by GPT-3 Brown et al. (2020) for text instruction generation, and Ip2p for image editing based on the instructions (See Supp.Sec.S2.3).
[Uncaptioned image] Figure 5: Human Psychophysics Experiment Results. The average proportion of images that human participants prefer our EmoEditor over other methods is 56%. Chance is 50% (red dotted line). Error bars are standard errors.    Method EMR(%)↑ ESR(%)↑ ENRD↓ ESS↓ CT 6.89 79.32 33.29 7.36 NST 34.42 92.01 34.45 18.57 Csty 11.51 85.52 41.47 36.64 Ip2p 2.53 67.76 9.39 12.71 LMS 11.51 77.38 26.13 19.74 w/o $\mathcal{P}$ 5.06 69.15 19.45 14.93 w/o $L_{emb}$ 43.35 91.62 22.53 16.00 Ours 50.20 92.86 23.98 16.27 Table 1: Quantitative Evaluation of Generated Images for All Competitive Methods and Ablated Models of our EmoEditor. See Sec.4 for the methods and evaluation metrics. See Sec.5.1 for the result analysis.
@@
Human Psychophysics Experiment. We evaluate the generated results of all methods on Amazon Mechanical Turk (MTurk)Turk (2012). We recruit 417 participants, with each participant undergoing 100 trials, yielding total 417,000 trials. All the experiments are conducted with the subjects’ informed consent and according to protocols approved by the Institutional Review Board of our institution. All participants are properly compensated. In each trial, participants engage in a “two-alternative forced choice" task. They are presented with two image stimuli and must select the one that most strongly evokes the target emotion in them. The two image stimuli consists of a result generated by our EmoEditor and a randomly sampled result from the five baselines introduced above. The trial presentation order and the binary choices for each trial are randomized. To assess the quality of data collection, we introduce control trials and use them as filtering criteria. We discard the data from those people who fail the control trials at least 17% of the time. See Supp.Sec.S2.4 for more details.
Evaluation Metrics. While CLIP Radford et al. (2021) is commonly used to assess text-image consistency, it doesn’t fully capture the complexity of emotions Widhoelzl, Takmaz (2024). Relying on CLIP for emotional alignment may overlook important nuances in images. Thus, we introduce four novel evaluation metrics to quantitatively assess emotion-evoked image generation performance.
Emotion Matching Ratio (EMR) and Emotion Similarity Ratio (ESR) assess the extent to which the generated images evoke target emotions. EMR: Similar to Goetschalckx et al. (2019), we also use our emotion predictor $\mathcal{P}$ to assess the emotional category of generated images and determine their alignment with target emotions. We acknowledge that using $\mathcal{P}$ as an evaluator may introduce favorable biases toward our model. However, the absence of existing emotion evaluators leaves us no choice but to rely on $\mathcal{P}$. ESR: $\mathcal{P}$ generates emotion probability distributions $e_{g}$ and $e_{s}$ for generated and source images. We compute ESR using Kullback-Leibler Divergence distance $KLD(\cdot)$, where $KLD(e_{g},e_{oh})<KLD(e_{s},e_{oh})$. $e_{oh}$ is the binary one-hot target emotion vector, indicating the impact of edits on generated images transitioning from the source emotion to the target emotion.
Emotion-Neutral Region Deviation (ENRD) and Edge Structure Similarity (ESS) assess the structural coherence and semantic consistency between source and generated images. ENRD: Using Grad-CAM Selvaraju et al. (2017) with $\mathcal{P}$, we binarize Grad-CAM maps at a 0.5 threshold to identify emotionally neutral regions on source images, valued at 0. Then, we compute the pixel-level L1 distance between these regions on source and generated images. ESS: Employing the Canny edge detection algorithm with thresholds of 200 and 500 on both source and generated images, we determine the L1 norm between their edge maps to quantify structural disparities.
@@ 5 Results
5.1 Quantitative Evaluation in Cross-Valence Scenarios
@@
Figure 5 shows human psychophysics experimental results (see Supp.Sec.S3.1 for more results). If all methods generate images that elicit target emotions equally effectively, human participants would make random choices, with 50% chance in the “two-alternative forced choice" tasks. Complimentary to psychophysics experiments, Table 1 presents quantitative results using four metrics from Sec.4. Observing Figure 5, we see that human participants consistently prefer the generated results of our EmoEditor over all competitive baselines, indicating its proficiency in evoking target emotions. This is supported by EmoEditor’s highest EMR and ESR scores in Table 1. Additionally, EmoEditor achieves moderately lower ENRD and ESS, suggesting maintained structural and semantic coherence in the generated images.
@@
Figure 6: Visualization of Generated Images from Different Methods. The target emotion is highlighted in red and the source image is framed with green. See Supp.Fig.S10 for more examples.
From Figure 5, we noted that our EmoEditor significantly outperforms Color Transfer (CT), indicating that emotion-evoked image generation involves more than global color and brightness adjustments. CT also has lower EMR and ESR than our EmoEditor in Table 1. As expected, CT only transfers colors, preserving the edges of the source image and hence, resulting in the lowest ESS scores. Similarly, our EmoEditor is significantly favored over Ip2p by human participants, indicating that current text-to-image diffusion models face challenges in understanding emotional cues and generating effective text instructions for editing emotion-evoked images. Despite Ip2p having comparable ENRD and ESS scores to our EmoEditor, its significantly lower EMR and ESR scores suggest that it makes minimal changes to source images, maintaining high structural and semantic similarity but failing to evoke target emotions. Additionally, our EmoEditor also outperforms Large Model Series (LMS) based on human preference scores in Figure 5, 39% in EMR, 15% in ESR, 2 in ENRD, and 3 in ESS in Table 1. This suggests that LMS accumulates biases and errors across all procedural steps, highlighting the importance of end-to-end training for emotion-evoked image generation.
Neural-Style-Transfer (NST) and CLIP-Styler (Csty) excel in inducing negative emotions through distortion and irregular textures on source images. They are preferred in positive-to-negative emotion manipulations, primarily in the upper quadrants of their confusion matrix (see Supp.Fig.S8). However, these methods render images challenging to comprehend, indicated by higher ENRD and ESS scores in Table 1, suggesting the distorted images lose structural similarities and semantic content. In contrast, our EmoEditor outperforms both methods in generating positively valenced images, with higher preference scores in the lower quadrants of the confusion matrix (see Supp.Fig.S8).
In ablation studies (Table 1), removing either the emotion predictor $\mathcal{P}$ (Row 6) or the alignment loss $\lambda L_{emb}$ (Row 7) significantly impairs performance. This leads to lower EMR and ESR scores, highlighting the crucial roles of both components in our EmoEditor for detecting emotional cues from source images, emulating human-like reasoning in generating editing instructions, and executing iterative refinements during inference. We also perform ablation studies on different max iterations for iterative emotion inference, showing that this iterative process improves the model’s performance, achieving stable and high-quality emotionally evocative images (see Supp.Sec.S3.3).
5.2 Visualization of Emotional Image Generation across Valence
Figure 1 and 6 present visualizations of generated images for all competitive methods in cross-valence scenarios (see Supp.Sec.S3.2 for more results). Color Transfer primarily replicates tonal characteristics from a reference image, often ineffective for significant emotional augmentation. Neural Style Transfer heavily relies on randomly selected reference images. While it produces artistic textures aligned with target emotions, it lacks in preserving essential semantic content in the source images. CLIPstyler excels in producing negative emotions but introduces challenging-to-interpret textures, as seen in Column 4. Ip2p demonstrates limited image generation abilities with single-word emotion prompts. For example, the fire in the source image remains almost identical to the generated image, due to the inability to comprehend emotions (Row 1, Column 5). Even after concatenating a series of large vision and language models before Ip2p, LMS does not always generate desirable outputs, triggering target emotions. This can be seen by the generated image of a smiling face, failing to evoke sadness in Row 2, Column 6.
In contrast, our EmoEditor needs only the target emotion and source image inputs, yet produces highly creative images that evoke target emotions while striving to maintain scene structures and semantic coherence as much as possible. For instance, in Row 1, Column 7, our method does not change the subject of the fire but turns the fire in the woods into a bonfire and generates a beautiful night sky in the background, thus evoking contentment. In Row 6, Column 7, it accurately adjusts the character’s facial expression while retaining the original content elements.
@@ 5.2 Visualization of Emotional Image Generation across Valence
Figure 7: Our EmoEditor can generalize to more challenging emotion editing scenarios. (a) Three examples are showcased, highlighting our EmoEditor’s ability to produce images evoking emotions of the same positive valence as the source images. (b) Two examples demonstrate how our EmoEditor can transform neutral real-world images to evoke either positive or negative emotions. Source images are framed in green. Target emotions are in red. See Sec.5.3 for the result analysis.
@@ 5.3 Generalization to Real-world Scenarios
Editing images within the same emotional valence presents greater challenges compared to cross-valence emotion image editing due to subtle differences in emotional cues. Figure 7(a) showcases our EmoEditor’s results within the same valence. Our EmoEditor displays a nuanced understanding of emotional cues and produces impressive creative outcomes. Example 1 transitions from awe to contentment, depicting a sunset scene with afterglow. Example 2 generates a blooming white flower with additional purple flowers to evoke excitement. Example 3 transforms a lively spring scene into a tranquil night scene with moonlight to evoke awe.
In real-world scenarios, some naturalistic images may evoke mixed emotions or no strong emotional responses at all. We challenge our EmoEditor to generate emotion-evoked images using neutral-valence source images randomly selected from the MSCOCO dataset Lin et al. (2014), commonly used for object detection. These images depict everyday life scenes with neutral emotions. Our EmoEditor applies positive and negative emotion transformations to these source images. Figure 7(b) illustrates the generated results. In Example 1, to inject amusement into an office space, our EmoEditor embellishes the walls with delightful watercolor paintings and infuses lively and vibrant colors into the room furniture. Conversely, to evoke fear, our model creates an aged desk and dimly lit walls. In Example 2, for an image featuring a courtyard with a teddy bear observing the garden, our method enriches the scene by adding a beautiful flower bed, creating an awe-inspiring effect. Conversely, to evoke feelings of disgust, our EmoEditor removes the flowers from the courtyard and transforms the view into a dilapidated scene with shabby clothes hanging on the walls and trash bags in the corner. It is crucial to emphasize that our EmoEditor is not trained in these neutral scenarios. All the generated results stem solely from the model’s own creativity and imagination, without any human-annotated editing instructions or scene interpretations.
@@ 6 Discussion
We present a new and important problem of emotion-evoked image generation. To address this, we present EmoEditor, an image diffusion model that understands emotional cues, creates implicit editing instructions aligned with human decisions, and manipulates image regions to evoke emotions, while maintaining coherent scene structures and semantics. Moreover, we contribute EmoPair dataset for model training. To benchmark all the methods, we introduce new evaluation metrics and establish standard protocols in emotion understanding, visualization, and reasoning. While our EmoEditor demonstrates superior quantitative performance and visually striking image generation results, we acknowledge several limitations in accurately handling fine-grained details of visual features on small faces within crowded scenes and generating emotion-evoked images without exacerbating semantic and structural disparities between source and target images. See Supp.Sec.S3.4 for details.
The problem of emotion-evoked image generation introduced here opens doors for research across AI, psychology, design, arts, and neuroscience. However, EmoEditor, like any technology, carries inherent risks. Misusing it, such as manipulating images to evoke undesirable emotions, could harm individuals’ mental well-being or mislead public feelings and opinions.
Supplementary Material
@@
In response to the absence of an image dataset specifically designed for emotion-evoke image generation, we introduce EmoPair, comprising two distinct subsets: the EmoPair-Annotated Subset (EPAS), encompassing 331,595 image pairs sourced from Ip2pBrooks et al. (2023) and annotated with emotion labels; and the EmoPair-Generated Subset (EPGS), featuring 6,949 pairs generated through text instructions specifying target emotions.
For the EmoPair-Generated Subset (EPGS), we formulated 50 general instructions that are agnostic to source images, prompting transitions to desired emotions across 8 categories using GPT-3 Brown et al. (2020). Human annotators then ranked these instructions based on efficacy within each emotion category, ultimately retaining the top ten. Figure S1 illustrates examples of these instructions.
We employ the following selection criteria to control the quality of generated image pairs of EPGS: (1) Using our emotion predictor $\mathcal{P}$, we analyze the generated images and only select those with a Top-1 classification confidence over 90% for the target emotions. (2) To ensure the preservation of similar scene structures, we utilize Structural Similarity Index (SSIM) Wang et al. (2004) and Learned Perceptual Image Patch Similarity (LPIPS) Zhang et al. (2018) for filtering the remaining outcomes. SSIM measures structural similarity between images, while LPIPS quantifies perceptual differences. Specifically, we require the generated images $\hat{x}$ and the source image $x$ to meet the conditions below: 0.3 < SSIM($x$, $\hat{x}$) < 0.6 and LPIPS($x$, $\hat{x}$) > 0.1. Ultimately, EPGS retains 6,949 image pairs.
To provide a more comprehensive overview of our dataset, additional image pair examples are presented in Figure S2 and S3.
@@
Our Emotion Encoder $\tau_{\theta}$ is a fully connected network designed for transforming one-hot encoded emotion input vectors into the structured emotion embedding $e$. The architecture of the model is composed of a series of fully connected layers that progressively increase the dimensionality of the input. Figure S4 shows the network structure of our emotion encoder $\tau_{\theta}$.
@@
S2.2 Implementation Details
Follow Wang et al. (2022), we initialize the weights of $\mathcal{E}$, $\mathcal{D}$, and $\epsilon_{\theta}$ with the pre-trained Ip2p Brooks et al. (2023) weights. Throughout the training process, we maintain the fixed parameters of $\mathcal{E}$ and $\mathcal{D}$, focusing on training $\tau_{\theta}$ and $\epsilon_{\theta}$. The frozen emotion predictor $\mathcal{P}$ is only used during inference. We conduct experiments with NVIDIA RTX A5000 GPUs, implementing our PyTorch-based framework using the ADAM optimizer. The max time step $T$ is set to 1,000.
S2.3 Baseline: Large Model Series
Emotion-evoked image generation involves three key steps: image understanding, instruction generation, and instruction-based image editing. To tackle these, we concatenate existing large language and vision models and introduce the baseline “Large Model Series" (LMS). This includes BLIP Li et al. (2022) for image captioning, followed by GPT-3 Brown et al. (2020) for text instruction generation, and Ip2p for image editing based on the instructions.
Figure S5 shows the workflow of LMS. First, we employ BLIP to comprehend the source image, generating an image caption corresponding to the source image. Subsequently, based on the generated caption, we utilize the sentence structure “There is an image with [Image caption]. To make me feel [Target Emotion], how to change the image? Give one instruction that starts with an action." to query GPT-3. Following this, GPT-3 generates an instruction to guide Ip2p in editing the source image, resulting in the final output.
Figure S5: Workflow of Large Model Series (LMS). We decompose emotion-evoked image generation into three steps: employing BLIP for image understanding, utilizing GPT-3 for instruction generation, and employing Ip2p for image editing. We refer to this approach of chaining multiple large models as the Large Model Series.
S2.4 Human Psychophysics Experiment
@@
(a) Each participant undergoes 6 randomly dispersed dummy trials within the real experiments. To assess if participants are making random selections, we use 6 image pairs with prominent emotional differences as references. Considering individual emotional response variability, participants are allowed a maximum of one incorrect choice in these dummy trials. Subjects exceeding an error rate of 1/6 are excluded, resulting in 11,400 trials. The outcomes of the 6 dummy trials are also excluded from the final statistics. Figure S6(c) provides an example of a dummy trial.

### v4

@@ 4 Experiment
Ideally, with source images and their classified emotions, our goal is to generate images evoking the other 7 target emotions. Considering emotional valence, we propose two tasks: within-valence emotion editing and cross-valence emotion editing. This division serves two purposes. Firstly, cross-valence emotion editing is more common in real-world scenarios, like enhancing positive emotions in stressful environments or creating fear in lively settings during Halloween. Secondly, images of the same valence may contain subtle differences challenging for humans to discern. For instance, an amusement park scene might evoke a mixture of excitement and amusement. We randomly select test images from EmoSet [33], ensuring they differ from those in our EmoPair dataset and that each source emotion category is evenly distributed. We then evaluate the competitive baselines and our EmoEditor on the cross-valence emotion editing task, with a total of 2,016 test images. We also present the results of applying our EmoEditor to more challenging emotion editing scenarios. To ensure a fair comparison, the input image size is standardized to 224$\times$224 across all methods.
@@
Baselines. We compare our EmoEditor with nine methods: (1) Color-Transfer (CT) [20]; (2) Neural-Style-Transfer (NST) [6]. Both methods require an additional reference image from the target emotion category as input. We randomly select these reference images from the EmoSet dataset. We also include representative text-driven image editing models (3) Csty [13]; (4) CtrlNet [38] (5) SDEdit [16]; (6) Ip2p [3] and (7) AIF [30]. We provide these models with the target emotion category as a one-word text prompt, such as “excitement". We also compare with (8) EmoEdit [34], which performs emotion-driven image editing by retrieving semantically similar examples from an emotion dictionary. In addition, we concatenate existing large language and vision models and introduce the baseline (9) “Large Model Series" (LMS). This includes GPT-4o [9] for image captioning, followed by GPT-o4 [18] for generating reasoning-based text instructions, and Ip2p for image editing based on the instructions (See Sec S2.2).
@@
Human Psychophysics Experiment. We evaluate the generated results of all methods on Amazon Mechanical Turk (MTurk) [26]. We recruit 136 participants, with each participant undergoing 180 trials, yielding a total of 24,480 trials. All the experiments are conducted with the subjects’ informed consent and according to protocols approved by the Institutional Review Board of our institution. All participants are properly compensated. In each trial, participants engage in a “two-alternative forced choice" task. They are presented with two image stimuli and must select the one that most strongly evokes the target emotion in them. The two image stimuli consist of a result generated by our EmoEditor and a randomly sampled result from the baselines introduced above. The trial presentation order and the binary choices for each trial are randomized. To assess the quality of data collection, we introduce control trials and use them as filtering criteria. We discard the data from those people who fail the control trials at least 17% of the time. See Sec S2.3 for more details.
Evaluation Metrics. While CLIP [21] assesses text-image consistency, it does not fully capture emotional complexity [31]. Thus, we introduce Emotional Structure Matching Index (ESMI): $ESMI=\alpha\times S_{str}+(1-\alpha)\times S_{emo}$ where $\alpha$ is a weighting coefficient and set to 0.5. ESMI evaluates emotion-evoked image generation from two aspects: evoking specific emotions and maintaining structural and semantic consistency between the source and generated images. Evoking Targeted Emotions. To assess how well the generated image expresses the target emotion, we use $\mathcal{P}$ to extract emotional distributions $e_{s}$ and $e_{g}$ from the source and generated images respectively. Next, we calculate the Kullback-Leibler Divergence between these distributions and the target emotion’s one-hot encoding $e_{oh}$, we get $D_{src}=KLD(e_{s},e_{oh})$ and $D_{gen}=KLD(e_{g},e_{oh})$. Ideally, the generated image should emotionally be closer to the target, that is, $D_{gen}<D_{src}$. To avoid the influence of outliers or extreme values, we define the raw emotional improvement as $\Delta D=\max(0,D_{src}-D_{gen})$. To further normalize this improvement relative to the original emotional distance, we compute the emotional evocation score: $S_{emo}=\Delta D/(D_{src}+\epsilon)$ where $\epsilon$ is a small constant to prevent division by zero. This score ranges from 0 to 1, with higher values indicating a greater emotional shift toward the target emotion relative to the source. Preserving Original Structures. We define two regions: the Emotion-Evoking Region ($R_{emo}$), which is responsible for triggering the source emotion (e.g. fire for anger) and is expected to be altered, and the Emotion-Neutral Region ($R_{neu}$), which should remain unchanged. We measure structural consistency by computing the pixel-level L1 differences between the source and generated images in these regions, denoted as $L_{emo}$ and $L_{neu}$. The structural preservation score is calculated as: $S_{str}=L_{emo}/(L_{emo}+L_{neu}+\epsilon)$ where $\epsilon$ is a small constant to avoid division by zero. This score reflects the proportion of changes in $R_{emo}$ relative to all changed pixels, with higher values indicating better preservation of neutral areas. We adopt two strategies to define $R_{emo}$: (1) CAM-based: $R_{emo}$ is obtained by applying Grad-CAM [24] to $\mathcal{P}$ and binarizing the resulting heatmap. (2) HA-based: $R_{emo}$ is derived by averaging manually annotated emotional regions of human annotators. See Sec S2.4 for more details.
ESMI(%) ↑ CT NST Csty CtrlNet SDEdit Ip2p AIF EmoEdit LMS Ours
CAM-based 28.33 47.48 34.33 32.94 35.91 27.37 29.71 32.94 35.40 51.56
HA-based 26.68 46.00 32.77 31.53 34.27 25.72 28.31 37.53 33.76 49.99
Table 1: Quantitative Evaluation of Generated Images for All Competitive Methods. The CAM-based metric computes ESMI using $R_{\text{emo}}$, derived by binarizing the Grad-CAM output from the Emo Predictor. The HA-based metric computes ESMI using $R_{\text{emo}}$, defined as the average of human-annotated regions. Best is in bold. Larger (↑) is better.
Figure 5: Results from Human Psychophysics Experiments. The proportions of images that human participants prefer our EmoEditor over all other methods on the x-axis are presented. Chance is 50% (red dotted line). Error bars are standard errors.
Variant Input of $\tau_{\theta}$ IEI $L_{emb}$ ESMI(%)↑
text $e_{oh}$ $e_{dir}$
1 5.85
2 50.24
3 7.36
4 8.51
5 47.36
6 49.80
7 (ours) 51.56
Table 2: Quantitative Evaluation in CAM-based ESMI for Model Variants of our EmoEditor. A tick indicates the inclusion of a component in the corresponding model variant, while a blank cell indicates its absence. The last row represents our full EmoEditor model. Best results are in bold. Larger (↑) is better.
@@ 5 Results
5.1 Evaluation in Cross-Valence Scenarios
@@
Quantitative Evaluation with SOTA. Figure 5 shows human psychophysics experimental results (see Sec S3.1 for more results). If all methods generate images that elicit target emotions equally effectively, human participants would make random choices, with a 50% chance in the “two-alternative forced choice" tasks. Observing Figure 5, we see that human participants consistently prefer the generated results of our EmoEditor over all competitive baselines, indicating its proficiency in evoking target emotions. Complimentary to psychophysics experiments, Table 1 presents quantitative results in CAM-based and HA-based ESMI, with our EmoEditor achieving the highest scores in both metrics. Notably, for each method, the scores remain consistent across the two evaluation metrics, indicating that the emotion predictor $\mathcal{P}$ identifies emotion-evoking regions in a manner aligned with human annotations. This consistency further supports the reliability and validity of the CAM-based ESMI metric.
Our EmoEditor significantly outperforms CT, demonstrating that emotion-evoked image generation requires more than global color adjustments. While NST and Csty can induce negative emotions through distortions and irregular textures, they often reduce image interpretability (see Fig S10). CtrlNet and SDEdit produce abstract textures that may evoke emotion but often fail to preserve the source image’s semantics. Human evaluations also prefer EmoEditor over Ip2p, underscoring the limitations of current diffusion models trained for general image generation when applied to emotion-driven tasks. Although both AIF and EmoEdit are designed for emotional editing, our EmoEditor achieves higher human preference and ESMI scores. AIF applies global filters, resulting in less targeted changes. EmoEdit maps image-emotion pairs into an embedding, but the limited and biased emotion dictionary leads to repetitive edits and poor performance on strongly emotional inputs. Finally, our EmoEditor outperforms LMS in both human preference and ESMI, suggesting that simply chaining large models may amplify biases and errors. Emotion-evoked generation demands a nuanced understanding of both the source image and target emotion, making end-to-end learning crucial.
@@
Figure 6: Visualization of Generated Images from Different Methods. Target emotion is highlighted in red and source image is framed in green. The blue number below each image is its CAM-based ESMI score, with higher values being better. More examples in Sec S3.2.
@@
Figure 7: Our EmoEditor can generalize to more challenging emotion editing scenarios. (a) Same-Valence Editing highlights our EmoEditor’s ability to produce images evoking emotions of the same positive valence as the source images. (b) Neutral-Valence Editing demonstrates how our EmoEditor can transform neutral real-world images to evoke positive or negative emotions. (c) Multi-Emotion Editing demonstrates its capability to generate images that evoke a wider range of emotions, providing more nuanced and diverse visual outputs. Source images are framed in green. Target emotions are in red.
Model Analysis. We conduct the model analysis and present the results in Table 2. In Row 1, the model variant takes only text as input and follows the structure of Ip2p. Although it is fine-tuned on our EmoPair dataset with the pre-trained CLIP model as $\tau_{\theta}$, the performance is the lowest among all variants, indicating that text guidance is insufficient. Adding IEI (Row 2) results in notable improvements over the version without IEI (Row 1), confirming its contribution to enhancing emotional consistency. Next, replacing the text input with a one-hot encoding of the target emotion $e_{oh}$ (Row 3) leads to a better result, highlighting the advantage of structured emotion encoding over text. Further improvement is observed when replacing $e_{oh}$ with an emotion editing direction $e_{dir}$ (Row 4), suggesting that incorporating information about the source emotion enables more precise and effective editing. Adding the alignment loss $L_{emb}$ (Row 6) improves performance over the variant without it (Row 5), confirming its complementary role in promoting emotionally consistent generation. Our final model (Row 7) integrates $e_{dir}$, IEI, and $L_{emb}$, enabling emotion-aware editing based on the source-target gap. This configuration achieves the highest score, demonstrating its superior emotional precision and generative diversity.
Visualization of Emotional Image Generation across Valence. Figure 1 and 6 present visualizations of generated images in cross-valence scenarios. CT primarily replicates tonal characteristics from a reference image, often failing to significantly enhance emotions. NST relies on randomly selected references, producing artistic textures aligned with target emotions but losing essential semantic content. AIF, dependent on emotional text descriptions, applies a global filter effect but struggles with localized editing. EmoEdit relies on fixed learned queries, leading to limited, context-incongruent edits (e.g., always adding balloons for amusement; Figure 1) and failing to edit local emotional cues such as fire flames or fireworks (Figure 6). Even with a richer vocabulary, such as LMS, which integrates three large models, it still struggles to generate desirable outputs. In contrast, our EmoEditor takes only a source image and target emotion as input, yet generates creative and emotionally aligned results while preserving structural and semantic coherence. By learning an emotion editing direction, it understands the source-target gap and applies precise edits. In Figure 6, where fire and fireworks dominate the original emotion, EmoEditor adapts its edits to the target: extinguishing flames and adding mountains to evoke excitement (a), or removing fireworks and darkening tones to convey sadness (b). In Figure 1, it replaces fire flames with lamps to elicit amusement. Notably, even when handling similar anger-inducing elements like fire, our EmoEditor produces diverse, context-aware edits tailored to the source image without any manual design, demonstrating a nuanced understanding of emotional cues. We also show CAM-based ESMI scores (in blue) for each result. Our EmoEditor achieves the highest scores, demonstrating its ability to accurately identify and effectively modify emotion-evoking regions in source images, striking a balance between evoking the target emotion and preserving the original structure.
@@ 5.2 Generalization to Real-world Scenarios
Editing images with the same emotional valence is more challenging due to subtle emotional cues. Figure 7a demonstrates EmoEditor’s ability to handle such tasks, as it removes a parachutist to reveal a breathtaking landscape, shifting excitement to awe. In real-world scenarios, some naturalistic images may evoke mixed emotions or no strong emotional responses at all. We challenge our EmoEditor to generate emotion-evoked images using neutral-valence source images randomly selected from the MSCOCO dataset [14]. These images depict scenes of everyday life with neutral emotions. EmoEditor applies positive and negative emotion transformations to these source images. It enhances the lushness of a forest to evoke contentment and transforms it into a barren landscape for disgust (Figure 7b). Notably, EmoEditor generates these results purely from its creativity, without any human-annotated editing instructions or scene interpretations.
Furthermore, EmoEditor can manage multi-emotion editing with the use of emotion encoding as input. We extend the single-target emotion scenario by randomly adding one additional emotion of the same valence as the target and then use EmoEditor for image generation. In Figure 7c, EmoEditor generates a majestic mountain in the flame area of the source image to evoke awe. After adding contentment, it enriches the area with dense forests.
@@ 6 Discussion
We introduce the challenging problem of emotion-evoked image generation and propose EmoEditor, an image diffusion model that understands emotional cues, creates implicit editing instructions aligned with human decisions, and manipulates image regions to evoke emotions, while maintaining coherent scene structures and semantics. Moreover, we contribute EmoPair dataset for model training. To benchmark all methods, we introduce a new evaluation metric that effectively measures the balance between evoking target emotions and maintaining the structural and semantic integrity of the source image.
While EmoEditor excels in quantitative performance and visual results, it struggles with fine details on faces (See Sec S3.4). While we adopt the 8-category Mikels model, real-world emotions are far more nuanced and complex. In particular, differentiating between positive emotions tends to be more challenging than distinguishing negative ones. Additionally, EmoEditor carries risks of misuse, such as evoking undesirable emotions, which could harm mental well-being or mislead public sentiment. All of these aspects warrant further investigation in future research.
@@
In response to the absence of an image dataset specifically designed for emotion-evoke image generation, we introduce EmoPair, comprising two distinct subsets: the EmoPair-Annotated Subset (EPAS), encompassing 331,595 image pairs sourced from Ip2p [3] and annotated with emotion labels; and the EmoPair-Generated Subset (EPGS), featuring 6,949 pairs generated through text instructions specifying target emotions.
For the EmoPair-Annotated Subset (EPAS), we use $\mathcal{P}$ to categorize the source and target images from the Ip2p dataset into eight different emotions. We remove samples where the emotional labels of the source and target images are consistent, then use the emotion labels to augment the Ip2p dataset, generating 331,595 image pairs.
@@
For the EmoPair-Generated Subset (EPGS), we formulated 50 general instructions that are agnostic to source images, prompting transitions to desired emotions across 8 categories using GPT-3 [4]. Human annotators then ranked these instructions based on efficacy within each emotion category, ultimately retaining the top ten. Figure S1 illustrates examples of these instructions.
We employ the following selection criteria to control the quality of generated image pairs of EPGS: (1) Using our emotion predictor $\mathcal{P}$, we analyze the generated images and only select those with a Top-1 classification confidence over 90% for the target emotions. (2) To ensure the preservation of similar scene structures, we utilize Structural Similarity Index (SSIM) [29] and Learned Perceptual Image Patch Similarity (LPIPS) [39] for filtering the remaining outcomes. SSIM measures structural similarity between images, while LPIPS quantifies perceptual differences. Specifically, we require the generated images $\hat{x}$ and the source image $x$ to meet the conditions below: $0.60.3<SSIM(x,\hat{x})<$ 0.6 and $LPIPS(x,\hat{x})>0.1$. Ultimately, EPGS retains 6,949 image pairs.
To ensure the quality of EPGS, we have human annotators re-annotate 300 images. The annotators view both the source and edited images, and we ask them to select the one that better evokes the target emotion. Since we need the target image to evoke the desired emotion more effectively than the source image, if the annotators choose the source image, we swap the source and target images in the original dataset to create a new subset. Each image is reviewed by three annotators, and the final subset is determined by majority vote. We then use this subset to fine-tune our EmoEditor.
To provide a comprehensive overview of our dataset, additional image pair examples are presented in Figure S3 and S3.
@@
Figure S4 shows the network structure of our emotion encoder $\tau_{\theta}$. We represent the target emotion as a one-hot encoding $e_{oh}$ and use the emotion predictor $\mathcal{P}$ to calculate the emotional state of the source image $e_{s}$. Then, we compute the emotion editing direction $e_{dir}$ by calculating the difference between $e_{oh}$ and $e_{s}$. Our Emotion Encoder $\tau_{\theta}$ is a fully connected network designed for transforming emotion editing direction $e_{dir}$ into the structured emotion embedding $e$. The architecture of the model is composed of a series of fully connected layers that progressively increase the dimensionality of the input.
@@
S2.2 Baseline: Large Model Series
Emotion-evoked image generation involves three key steps: image understanding, instruction generation, and instruction-based image editing. To tackle these, we concatenate existing large language and vision models and introduce the baseline “Large Model Series" (LMS). This includes GPT-4o [9] for image captioning, followed by GPT-o4 [18] for generating reasoning-based text instructions, and Ip2p for image editing based on the instructions.
Figure S5: Workflow of Large Model Series (LMS). We decompose emotion-evoked image generation into three steps: employing GPT-4o for image understanding, utilizing GPT-o4 for instruction generation, and employing Ip2p for image editing. We refer to this approach of chaining multiple large models as the Large Model Series.
Figure S5 shows the workflow of LMS. First, we employ GPT-4o to comprehend the source image, generating an image caption corresponding to the source image. Subsequently, based on the generated caption, we utilize the sentence structure “There is an image with [Image caption]. To make me feel [Target Emotion], how to change the image? Give one instruction that starts with an action." to query GPT-o4. Following this, GPT-o4 generates an instruction to guide Ip2p in editing the source image, resulting in the final output.
S2.3 Human Psychophysics Experiment
@@
(a) Each participant undergoes 6 randomly dispersed dummy trials within the real experiments. To assess if participants are making random selections, we use 6 image pairs with prominent emotional differences as references. Considering individual emotional response variability, participants are allowed a maximum of one incorrect choice in these dummy trials. Subjects exceeding an error rate of 1/6 are excluded, resulting in 24,480 trials. The outcomes of the 6 dummy trials are also excluded from the final result analysis. Figure S6(c) provides an example of a dummy trial.